OpenStudy (anonymous):

Describe how the graph of y= x2 can be transformed to the graph of the given equation. y = (x - 8)2 - 12 Shift the graph of y = x2 left 8 units and then down 12 units. Shift the graph of y = x2 right 8 units and then up 12 units. Shift the graph of y = x2 down 8 units and then left 12 units. Shift the graph of y = x2 right 8 units and then down 12 units.

OpenStudy (anonymous):

y = (x - 8)^2 - 12 is like: y = (x-h)^2 + k where h, and k represent the x and y value of the vertex of the parabola. if h = 8, then what does k=?

OpenStudy (anonymous):

k=12

OpenStudy (anonymous):

@DemolisionWolf

OpenStudy (anonymous):

pretty much, i'll just add the negative to it, so its y = -12

OpenStudy (anonymous):

haha sorry, k=-12, not y

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

so it would be down by 12 correct?

OpenStudy (anonymous):

so the value of k =-12 means the vertext is 12 down from the orgin.

OpenStudy (anonymous):

and the h value of 8 means the vertex is 8 units to the right of the orgin

OpenStudy (anonymous):

makes ense?

OpenStudy (anonymous):

yes (: thank you

OpenStudy (anonymous):

^_^